I need help solving these kind of problems........through (-4,0); parallel to y=3/4x-2. Idk how to begin it.
Answer by oscargut(2103)

Since the required line is parallel it has the same slope (3/4)
y = (3/4)x +b
Since (-4,0) is in the line
0 = (3/4)(-4) +b
0 = -3+b
b = 3
Answer: y = (3/4)x+3
You can ask me more at: mthman@gmail.com
Thanks